OpenStudy (anonymous):

whats is the answer ƒ(x ) = 3x + 1 and ƒ -1 = , then ƒ(2) = 1 7 1/3

OpenStudy (shamim):

f(x)=3x+1 f(2)=3*2+1=?

OpenStudy (shamim):

Just x=2

OpenStudy (anonymous):

If ƒ(a) = K, then which of the following must be true? ƒ(K) = a ƒ -1(K) = a ƒ -1(a) = K

OpenStudy (anonymous):

what is the answer